### v2

# Physics Context Builders: A Modular Framework for Physical Reasoning in Vision-Language Models

Abstract
Physical reasoning, which involves interpreting object behaviors within dynamic environments, remains a significant challenge for Vision-Language Models (VLMs). The limitations in physical reasoning arise from an inability to translate learned knowledge into predictions about physical behavior. We perform a careful study to show how continual fine-tuning can mitigate this issue. However, fine-tuning is expensive for large models and impractical to repeatedly perform for every task. This necessitates the creation of modular and scalable ways to teach VLMs about physical reasoning. To that end, we introduce Physics Context Builders (PCBs), a novel modular framework where specialized VLMs are fine-tuned to generate detailed physical scene descriptions. These can be used as physical contexts for larger VLMs to enhance their reasoning capabilities. PCBs enable the separation of visual perception from reasoning, allowing us to analyze their relative contributions to physical understanding. We perform careful experiments on CLEVRER and on Falling Tower, a stability detection dataset with both simulated and real-world scenes, to demonstrate that PCBs provide substantial performance improvements, increasing average accuracy by up to 13.8% on complex physical reasoning tasks. Notably, PCBs show strong Sim2Real transfer, successfully generalizing from simulated training data to real-world scenes. Our work demonstrates that enhancing visual perception through modular, simulation-trained components offers a practical approach to improving physical reasoning in VLMs, while providing insights into the factors affecting physical understanding in these models.
1 Introduction
Physical reasoning is a fundamental component of human intelligence, enabling interpretation of complex interactions, prediction of future events, and understanding of causal relationships in real-world environments (Kubricht et al.,, 2017). For humans, the ability to understand the physical world is developed early and operates intuitively (McCloskey et al.,, 1983; Baillargeon,, 2004; Téglás et al.,, 2011). However, physical reasoning remains a significant challenge for artificial intelligence (AI) systems (Lake et al.,, 2017; Srivastava et al.,, 2022; Yi et al.,, 2019; Chow et al.,, 2025), despite advances in computer vision and Vision-Language Models (VLMs) (Radford et al.,, 2021; Alayrac et al.,, 2022; Li et al.,, 2023; Zhang et al.,, 2024; Sun et al.,, 2024; Li et al.,, 2024; Ventura et al.,, 2025),
VLMs are remarkably successful on many predictive problems, spanning a broad swathe of tasks. This makes their use central to real-world applications requiring expertise in a variety of tasks such as robotics and embodied AI (Liu et al.,, 2024; Guo et al.,, 2024). However, current VLMs consistently fail at physical reasoning tasks, struggling with basic spatial relationships (e.g., object positioning, counting) (Rahmanzadehgervi et al.,, 2024), object attributes (Wang et al.,, 2023), and physical interactions (e.g., stability assessment, dynamics prediction) (Ghaffari and Krishnaswamy,, 2024; Chow et al.,, 2025). While humans leverage causal or physically guided knowledge for physical understanding of the world (Battaglia et al.,, 2013), the mechanisms by which VLMs make predictions are not well understood. This leads to two important questions: what factors contribute to the lack of physical understanding in VLMs and how to improve it?
One potential explanation for such limitations lies in VLMs’ training data. VLMs are created by fusing the representations of image and text encoders using datasets like MSCOCO (Lin et al.,, 2014) and Conceptual Captions (Sharma et al.,, 2018), which focus on general scene descriptions but lack annotations of physical relationships. Our experiments with standard benchmarks like CLEVRER (Yi et al.,, 2019) confirm this hypothesis, demonstrating that fine-tuning on physics-focused data can enable a relatively small VLM to approach state-of-the-art results achieved by specialized architectures (Yi et al.,, 2019; Ding et al., 2021a, ). However, as the ecosystem of open and closed source models grows, repeatedly fine-tuning each VLM for every regime of physical concept it cannot reason about becomes impractical; there is a need for practical, performant and modular tools to augment VLM capabilities.
To this end, we introduce Physics Context Builders (PCBs); these are specialized VLMs that are fine-tuned on simulation data to generate fine-grained physical descriptions that can be used by larger VLMs. Simulation can provide precise annotations of controllable physical interactions of interest, which can serve as scalable training data to teach models about physical phenomena. We show how to amortize the knowledge of physical concepts into PCBs and how they can transfer this knowledge to larger VLMs; this enables us to effectively separate visual perception from reasoning, using PCBs as perception modules, while leaving reasoning to larger models. See Figure 1 for a demonstration of PCBs.
Figure 1: Physics Context Builders (PCBs) pipeline. The training phase (top) shows how physics simulators generate images/videos with corresponding annotations, which are converted into two types of physical descriptions: Human-like Narration (providing natural language scene descriptions with object properties and spatial relationships) and Structured Physics (offering frame-by-frame structured descriptions with physical properties). These descriptions serve as training data for fine-tuning a relatively small VLM into specialized PCBs. During the inference phase (bottom), a trained PCB processes a new image/video and generates detailed physical context about the scene, which is then provided to a foundation model (e.g., GPT-4o) alongside a user query to produce physically grounded responses.
PCBs offer a practical and modular approach to enhance physical reasoning of existing large-scale VLMs without requiring expensive or sometimes infeasible modification of the larger foundation models. Our experiments on standard benchmarks demonstrate effective integration of PCBs with standard commercial VLMs. Moreover, they show that PCBs trained with simulation data successfully generalize to real-world scenarios, effectively performing simulation-to-reality (Sim2Real) transfer. In summary, our contributions are as follows:
  1. 1.
    We introduce a modular framework where specialized VLMs are fine-tuned to generate detailed physical scene descriptions, enhancing the physical reasoning capabilities of foundation models without their modification.
  2. 2.
    We demonstrate the effectiveness of separating visual perception from reasoning through our PCB approach, providing insights into how each component contributes to physical understanding.
  3. 3.
    We show how simulation data can be leveraged to train specialized modules that successfully transfer to real-world scenarios, avoiding the need for expensive simulation during inference.
2 Related Work
Physical Reasoning in Vision Models
Physical reasoning presents significant challenges for vision models, including both specialized architectures and large-scale VLMs. Recent benchmarks demonstrate that current VLMs struggle with basic physics understanding, often failing at tasks such as counting, depth reasoning, and physical interaction prediction (Ghaffari and Krishnaswamy,, 2024; Li et al.,, 2024; Ventura et al.,, 2025). These limitations are not unique to VLMs; physical reasoning has also proven difficult for single-purpose models, necessitating specialized architectures such as physics-inspired predictive models (Battaglia et al.,, 2016; Guen and Thome,, 2020; Duan et al.,, 2022), neural-symbolic executors (Chen et al.,, 2021; Ishay et al.,, 2024; Zheng et al.,, 2024), differentiable physics engines (Wu et al.,, 2015; Ding et al., 2021b, ), simulation-in-the-loop approaches (Liu et al.,, 2022; Zhu et al.,, 2025), and task-specific architectures (Ding et al., 2021a, ). However, large VLMs are not amenable to easy modifications of their neural network architectures and require alternative approaches to enhance their capabilities, which we consider in this work.
Evaluating Physical Reasoning in VLMs
Several recent works have evaluated physical reasoning capabilities in vision-language models. Nagar et al., (2024) benchmark zero-shot visual reasoning in both large language models (LLMs) and VLMs. They find that underlying LLMs, when provided with textual scene descriptions, consistently outperform VLMs that use visual embeddings. Their analysis shows that this performance gap is due to VLMs’ difficulty in translating visual information into accurate representations for reasoning. Our work builds on this insight by developing a modular approach to bridge this gap without requiring extensive retraining of VLMs.
Concurrently, Chow et al., (2025) introduced PhysBench, a comprehensive benchmark for evaluating physical understanding in VLMs. Their work reveals that VLMs’ physical reasoning do not scale proportionally with model size, training data, or input frame count. They identify perceptual and knowledge gaps as the primary sources of errors and propose using vision foundation models like Depth Anything (Yang et al.,, 2024) and SAM (Kirillov et al.,, 2023) to enhance visual perception. While our approach shares the goal of improving physical reasoning, we take a different direction by leveraging simulation data to train specialized context builders rather than relying on generic vision foundation models.
Simulation for Physical Understanding
Simulation has long been recognized as a valuable tool for teaching machines about physical dynamics (Battaglia et al.,, 2013). Previous approaches have incorporated simulation directly into the inference pipeline (Wu et al.,, 2015, 2017; Liu et al.,, 2022), requiring computationally expensive simulators during inference. In contrast, our method leverages simulation only during the training phase to generate rich physical descriptions, eliminating the need for simulation during inference while still benefiting from the detailed annotations that simulations provide.
Enhancing General Capabilities of VLMs
Recent work has explored modular approaches to enhance VLM capabilities without full model retraining. Chain-of-thought prompting techniques (Wei et al.,, 2022) leverage large language models’ reasoning abilities by encouraging step-by-step thinking, while multimodal chain-of-thought approaches (Zhang et al.,, 2023) extend this to vision-language tasks. Our Physics Context Builders build on these ideas by creating modular components that specialize in translating visual inputs into detailed physical descriptions.
3 Understanding the Effect of Training Data on Physical Reasoning
Here, we investigate why current VLMs struggle with physical reasoning tasks. We first describe our experimental setup, including datasets, models, and training procedures. We then present results demonstrating how fine-tuning with physics-focused data significantly improves performance of VLMs. Finally, we explore the nuanced influence of training data quantity vs. quality.
3.1 Experimental Setup
We utilize two benchmarks: CLEVRER (Yi et al.,, 2019) for dynamic physical reasoning and Falling Tower for static stability detection.
CLEVRER
is an established benchmark containing 10,000 training videos, 5,000 validation videos, and 5,000 test videos, each paired with multiple questions across four categories: descriptive, explanatory, predictive, and counterfactual. The training set consists of 109,952 descriptive, 16,799 explanatory, 7,179 predictive, and 18,642 counterfactual questions. All descriptive questions are open-ended, while other types are all multi-choice questions with at most four options, where there can be multiple correct options. We report the accuracy of the models stratified by the question types. For multi-choice questions, we report both per-option accuracy, which measures the model’s overall correctness on single options across all questions, and per-question accuracy that measures the model’s performance on full questions, requiring all the choices to be selected correctly. Since the test set answers are not publicly available, we evaluate on the validation set. The dataset’s questions are derived from simulation, making it particularly suitable for our investigation into how simulation-based training data affects physical reasoning.
Falling Tower
Figure 2: Datasets used for experiments.
is a dataset we create to complement CLEVRER by focusing on static physics. Similar to ShapeStacks (Groth et al.,, 2018) and block towers in (Lerer et al.,, 2016), it features stacked objects but extends previous work by including question-answer pairs and simulation-generated annotations. The dataset contains 4,864 images of stacked objects (with 15 object types across 3 shapes and 5 colors) and 72,775 QA pairs, with 75% used for training and 25% for evaluation. Questions and corresponding answers are generated by applying transformation functions to the simulation annotations, converting it into natural language questions and calculating the answers. For instance, to assess the stability of a stacked tower of objects, we compare the initial and final positions of the objects to determine if significant movement occurred. If an object remains within a defined threshold, we conclude it has remained stable. Questions are then categorized into descriptive (e.g., ”How many objects are in the scene?”) and stability (e.g., ”Will this collection of objects remain stationary?”). In addition to the simulated data, the dataset enables the evaluation of Sim2Real transfer; we include 20 real-world images captured with 3D-printed objects with 100 human-generated QAs. An example of the real-world images can be seen in Figure 2. For more details on Falling Tower, see Appendix C.
Models and evaluation.
We evaluate several zero-shot baselines, including GPT-4o, GPT-4o-mini (Achiam et al.,, 2023; Hurst et al.,, 2024), Gemini-1.5-Pro (Team et al.,, 2024), and PaliGemma-3B-mix (a variant fine-tuned on academic datasets) (Beyer et al.,, 2024) with chain-of-thought prompting to encourage explicit reasoning. To study training data effects, we fine-tune PaliGemma-3B on CLEVRER and Falling Tower independently using Low-Rank Adaptation (LoRA) (Hu et al.,, 2021), minimizing the auto-regressive negative log-likelihood of the answers, conditioned on the questions and input video/image. For videos, we sample 8 frames and append it to the input context, except for the Gemini model, where we use the entire video as the input since it supports video input. For all models, we take the final answer after the reasoning chain for evaluation. See section A.1 for more details on the training procedure.
Framing the questions.
For all questions, including the open-ended and multi-choice ones, we provide the potential options in the question statement. Moreover, for CLEVRER, we frame the multi-choice questions as multiple binary questions, where we ask if each choice can be an answer to the question or not. This yields significant improvement in the accuracy of all the evaluated models, as reported in section B.1.
3.2 Is Physics-Focused Training Data Enough?
Table 1: Performance of zero-shot VLMs and the fine-tuned model on the Falling Tower benchmark. The second number after the first slash is the Sim2Real accuracy.
Category Model Descriptive [sim acc. / real acc.] Stability [sim acc. / real acc. / real F1] num. obj. num. $\uparrow\downarrow↑ ↓ obj. absent\uparrow\downarrow$↑ ↓ obj. stable tower stable Random Random 33.3 25.0 14.3 50.0 50.0 Zero-shot CoT GPT-4o 99.3 / 100.0 91.4 / 100.0 99.4 / 95.0 56.9 / 60.0 / 63.3 59.6 / 55.0 / 54.2 GPT-4o-mini 94.9 / 89.5 61.3 / 84.2 87.9 / 73.7 49.0 / 52.6 / 55.0 53.1 / 36.8 / 19.8 Gemini 1.5 Pro 97.2 / 95.0 89.9 / 100.0 97.8 / 95.0 54.6 / 80.0 / 80.0 60.5 / 60.0 / 60.1 PaliGemma-3B-mix 91.4 / 90.0 44.5 / 70.0 73.0 / 65.0 51.0 / 75.0 / 68.6 39.1 / 65.0 / 51.2 Fine-tuned QA PaliGemma-3B-Fine-Tuned 100.0 / 100.0 100.0 / 100.0 100.0 / 95.0 84.6 / 70.0 / 73.0 87.6 / 65.0 / 64.4
Table 2: Performance of zero-shot VLMs and the fine-tuned model on the CLEVRER benchmark.
Category Model Descriptive Explanatory Predictive Counterfactual per ques. per opt. per ques. per opt. per ques. per opt. Random Random 28.8 11.8 50.0 25.0 50.0 7.4 50.0 Zero-shot CoT GPT-4o 62.7 30.7 65.5 30.3 47.6 18.7 60.2 GPT-4o-mini 49.5 9.3 51.8 44.9 45.1 15.6 51.0 Gemini 1.5 Pro 58.6 15.7 61.2 32.0 49.6 17.6 55.6 PaliGemma-3B-mix 38.9 6.6 33.4 44.7 48.7 7.7 49.8 Fine-tuned QA PaliGemma-3B-Fine-Tuned 92.9 94.7 98.2 83.6 83.6 68.4 88.7
Tables 1 and 2 present our results for the Falling Tower and CLEVRER benchmarks, respectively. Several key insights emerge from these results:
Zero-shot VLMs struggle with physical reasoning despite strong descriptive capabilities.
On Falling Tower, large models like GPT-4o and Gemini-1.5-Pro demonstrate near-perfect accuracy (95-100%) on descriptive questions, both in simulated and real environments. However, their performance drops substantially on stability questions, with accuracy ranging from only 55-60% (barely above random guessing) on stability detection. Similarly, on CLEVRER, while GPT-4o achieves 62.7% accuracy on descriptive questions, its performance falls to 30.7%, 30.3%, and 18.7% on explanatory, predictive, and counterfactual questions, respectively. This substantial performance gap highlights that while VLMs have developed strong capabilities for high-level scene understanding and description, they lack the abilities needed to predict outcomes or explain causal relationships accurately.
Fine-tuning with physics-focused data substantially improves physical reasoning.
Unsurprisingly, the fine-tuned PaliGemma-3B model shows dramatic improvements across all tasks. On Falling Tower, it achieves perfect descriptive accuracy (100%) and substantial gains in stability prediction (up to 28% improvement compared to GPT-4o). On CLEVRER, the improvements are even more significant, reaching close to the state-of-the-art results by Aloe, a specialized architecture trained on CLEVRER (Descriptive: 94.0 %, Explanatory: 96.0%, Predictive: 87.5%, Counterfactual: 75.6%) (Ding et al., 2021a, ). These results demonstrate that targeted fine-tuning with physics-focused data can significantly enhance physical reasoning capabilities, enabling a relatively small 3B parameter model to outperform much larger state-of-the-art models.
Sim2Real transfer is successful.
The fine-tuned model maintains strong performance when transferring from simulated to real-world images in Falling Tower. It achieves 100% accuracy on descriptive tasks and 70.0% / 65.0% accuracy on object / tower stability questions, respectively.
3.3 Ablation Studies: Not All Data is Equal
To gain deeper insights into how training data composition affects physical reasoning capabilities, we conduct several ablation studies. We provide more ablations in Appendix B.
Table 3: Effect of including different QA types in fine-tuning PaliGemma-3B on Falling Tower.
Fine-tuning Data Descriptive [acc.] Stability [acc.] num. obj. num. $\uparrow\downarrow↑ ↓ obj. absent\uparrow\downarrow$↑ ↓ obj. stable tower stable No Fine-tuning 50.9 21.2 8.5 51.0 39.1 Stability QAs 52.4 26.0 41.2 84.4 86.5 Descriptive QAs 100.0 100.0 100.0 51.0 39.1 All QAs 100.0 100.0 100.0 84.6 87.6
Figure 3: The amount of training data vs. performance on different QA types in Falling Tower.
Specificity of QA types in fine-tuning.
Table 3 shows the impact of using different subsets of QA pairs for fine-tuning on Falling Tower. Fine-tuning with only descriptive questions achieves perfect descriptive accuracy but fails to improve stability prediction (remaining at baseline levels of 51.0% and 39.1%). Interestingly, fine-tuning with only stability questions not only improves stability prediction substantially but also yields moderate improvements in descriptive tasks, particularly for object identification (41.2% vs. 8.5% baseline). However, the best performance comes from fine-tuning with the complete QA dataset. This highlights the need for carefully selecting the fine-tuning data to cover all potential QA types, which can be seen as a limitation of targeted fine-tuning in VLMs.
Data efficiency varies by task type.
[…]
4 Physics Context Builders: A Modular Approach to Physical Reasoning
While fine-tuning with physics-focused data can significantly improve physical reasoning in VLMs, this approach faces practical limitations. Fine-tuning state-of-the-art foundation models like GPT-4o or Gemini is often expensive or even impossible due to their closed-source nature and computational requirements. Furthermore, the task-specific nature of fine-tuning, as seen in section 3, means separate models may be needed for different physical reasoning tasks, which limits the generalizability of the approach. To address these challenges, we introduce Physics Context Builders (PCBs) - a modular and efficient approach that enhances physical reasoning capabilities of existing VLMs without modifying them directly. PCBs leverage the strong in-context learning abilities of large language models (Brown,, 2020; Wei et al.,, 2022), which have shown impressive performance with textual scene descriptions (Nagar et al.,, 2024).
4.1 Methodology
PCBs are specialized VLMs fine-tuned on simulation data to generate detailed physical descriptions of visual scenes. Rather than directly answering questions, PCBs act as perception modules that translate visual inputs into rich textual descriptions capturing the physical properties and dynamics of a scene. These descriptions then serve as enhanced context for larger VLMs, enabling more accurate physical reasoning through in-context learning. Figure 1 provides an overview of our approach. PCBs offer several advantages:
  • Modularity: PCBs can be fine-tuned and deployed independently of the foundation model.
  • Efficiency: Only smaller, specialized models need to be fine-tuned, rather than larger VLMs.
  • Flexibility: Different PCBs can be developed for different physical phenomena, creating a toolbox of physical reasoning enhancers.
  • Compatibility: PCBs work with any (vision) language model capable of in-context learning, including closed-source commercial models.
4.1.1 Physical Context Generation
Using annotations from the simulator, PCBs generate question-agnostic descriptions that capture the physical essence of a scene. We consider two approaches for context generation: (1) Human-like Narration (HN) that uses natural language descriptions of the scene’s physical properties and events that can be more intuitive for foundation models, and (2) Structured Physics (SP), which provides frame-by-frame structured observations in a format similar to a physics simulation output, with standardized tags for physical properties. This structured approach helps models capture precise temporal relationships and physical properties. For the stability detection task in Falling Tower, we generate descriptions, following the human-like narration, including object types, locations, and stacking arrangements. Figure 4 provides examples for these description formats.
Figure 4: PCB generation formats. (left) Human-like narration for CLEVRER. (top right) Structured physics for CLEVRER. (bottom right) Human-like narration for Falling Tower.
4.1.2 Training PCBs
We train PCBs by fine-tuning a pre-trained VLM to generate physical descriptions from visual inputs. Specifically:
  1. 1.
    We use PaliGemma-3B as the base pre-trained VLM due to its strong vision-language capabilities and reasonable computational requirements.
  2. 2.
    Similar to section 3, we apply LoRA-based fine-tuning (Hu et al.,, 2021), minimizing the auto-regressive negative log-likelihood of the context descriptions conditioned on the input video/image. See section A.2 for more details.
  3. 3.
    For videos (e.g., in CLEVRER), we sample 8 frames and append them to the input context.
  4. 4.
    The fine-tuning process targets both vision and language components, ensuring proper alignment between visual features and linguistic representations.
The training data is generated from the same simulations used to create the QA pairs, but focuses on comprehensive scene descriptions rather than specific questions and answers. This approach enables PCBs to provide rich context independent of the specific reasoning task.
4.2 Experimental Results
We evaluate the effectiveness of PCBs by integrating them with three foundation models: GPT-4o, GPT-4o-mini, and Gemini-1.5-Pro. In each experiment, we first pass the visual input to the appropriate PCB, which generates a physical description. This description is then provided as additional context to the foundation model along with the user’s question and the video/image.
Table 4: Performance of VLMs augmented with Physics Context Builders (PCBs) on the Falling Tower benchmark. HN refers to Human Narration style PCB. The numbers in parentheses indicate improvements over the respective zero-shot baselines. The second number after the slash is the Sim2Real accuracy and the third number after the second slash is the F1 score on Sim2Real.
Category Model Descriptive [sim acc. / real acc.] Stability [sim acc. / real acc. / real F1] num. obj. num. $\uparrow\downarrow↑ ↓ obj. absent\uparrow\downarrow$↑ ↓ obj. stable tower stable Zero-shot CoT GPT-4o 99.3 / 100.0 91.4 / 100.0 99.4 / 95.0 56.9 / 60.0 / 63.3 59.6 / 55.0 / 54.2 GPT-4o-mini 94.9 / 89.5 61.3 / 84.2 87.9 / 73.7 49.0 / 52.6 / 55.0 53.1 / 36.8 / 19.8 Gemini 1.5 Pro 97.2 / 95.0 89.9 / 100.0 97.8 / 95.0 54.6 / 80.0 / 80.0 60.5 / 60.0 / 60.1 VLM + PCB (HN) GPT-4o-PCB 99.5 / 100.0 97.6 / 100.0 99.5 / 95.0 76.7 / 75.0 / 73.8 85.1 / 70.0 / 65.6 (+0.2) / (0.0) (+6.2) / (0.0) (+0.1) / (0.0) (+19.8) / (+15.0) / (+10.5) (+25.5) / (+15.0) / (+11.4) GPT-4o-mini-PCB 99.9 / 95.0 74.3 / 90.0 97.5 / 95.0 75.0 / 70.0 / 73.0 84.7 / 40.0 / 33.8 (+5.0) / (+5.5) (+13.0) / (+5.8) (+9.6) / (+21.3) (+26.0) / (+17.4) / (+18.0) (+31.6) / (+3.2) / (+14.0) Gemini 1.5 Pro-PCB 97.9 / 100.0 97.5 / 100.0 97.4 / 94.7 75.9 / 73.7 / 76.7 84.9 / 57.9 / 59.1 (+0.7) / (+5.0) (+7.6) / (0.0) (-0.4) / (-0.3) (+21.3) / (-6.3) / (-3.3) (+24.4) / (-2.1) / (-1.0)
Table 5: Performance of VLMs augmented with Physics Context Builders (PCBs) on the CLEVRER benchmark. HN is Human Narration and SP is Structured Physics. The numbers in parentheses indicate improvements over the respective zero-shot baselines.
Category Model Descriptive Explanatory Counterfactual per ques. per opt. per ques. per opt. Zero-shot CoT GPT-4o 62.7 30.7 65.5 18.7 60.2 GPT-4o-mini 49.5 9.3 51.8 15.6 51.0 Gemini 1.5 Pro 58.6 15.7 61.2 17.6 55.6 VLM + PCB (HN) GPT-4o-PCB 75.6 (+12.9) 41.6 (+10.9) 67.0 (+1.5) 28.2 (+9.5) 68.4 (+8.2) GPT-4o-mini-PCB 65.7 (+16.2) 26.8 (+17.5) 62.2 (+10.4) 17.3 (+1.7) 52.8 (+1.8) Gemini 1.5 Pro-PCB 72.8 (+14.2) 35.6 (+19.9) 70.8 (+9.6) 26.2 (+8.6) 64.9 (+9.3) VLM + PCB (SP) GPT-4o-PCB 70.0 (+7.3) 34.9 (+4.2) 61.1 (-4.4) 19.2 (+0.5) 63.3 (+3.1) GPT-4o-mini-PCB 58.6 (+9.1) 19.2 (+9.9) 58.9 (+7.1) 16.2 (+0.6) 51.8 (+0.8) Gemini 1.5 Pro-PCB 67.4 (+8.8) 30.5 (+14.8) 68.7 (+7.5) 21.1 (+3.5) 60.5 (+4.9)
4.2.1 PCB Performance on Falling Tower
Table 4 shows the performance of foundation models augmented with PCBs on the Falling Tower benchmark:
  • Substantial improvement in stability tasks: PCBs provide remarkable gains in stability prediction, with GPT-4o-PCB showing up to 25.5% improvement in tower stability assessment. These gains are even more pronounced for smaller models like GPT-4o-mini, which sees a 31.6% improvement in tower stability prediction.
  • Modest gains in descriptive tasks: Since foundation models already perform well in descriptive tasks, PCBs offer limited additional benefits in this area, with improvements mainly in the more challenging descriptive tasks like identifying objects above/below others (e.g., 13.0% improvement for GPT-4o-mini).
  • Effective Sim2Real transfer: PCB-augmented models show improved generalization to real-world scenarios, with GPT-4o-PCB achieving a 15.0% gain in real-world tower stability prediction compared to zero-shot.
  • Model size effects: Interestingly, the smaller GPT-4o-mini model shows greater relative improvements when augmented with PCBs compared to larger ones, possibly due to its more limited perception.
4.2.2 PCB Performance on CLEVRER
Table 5 presents results for the more challenging CLEVRER benchmark, which requires reasoning about dynamic physics in videos:
  • Human Narration vs. Structured Physics: Human-like narration (HN) consistently outperforms structured physics (SP) descriptions across all models and question types. This can be due to foundation models’ better understanding of natural language descriptions compared to more structured, technical formats.
  • Strong improvements in descriptive and explanatory tasks: PCBs provide substantial gains in descriptive accuracy (up to 16.2% for GPT-4o-mini) and explanatory reasoning (up to 19.9% for Gemini 1.5 Pro).
  • Limited gains in counterfactual reasoning: While PCBs improve counterfactual reasoning, the gains are more modest (around 1.7-9.5%), highlighting the inherent difficulty of this task even with enhanced context.
Remark. For CLEVRER, we omit PCB evaluation on predictive questions since our current PCBs are designed to describe observed scenes rather than predict future events. This limitation could be addressed in future work by developing PCBs that can describe potential future states, which can be easily obtained using simulation.
4.3 Multi-Agent Framework for PCB Integration
Figure 5: Multi-agent framework for PCB integration. A triage agent selects the appropriate PCB based on the user query, and the PCB generates a detailed scene description that enriches the context of the foundation model.
Table 6: Performance metrics for selecting the correct context builder agent in a multi-agent framework (two agents).
Task Metric GPT-4o / GPT-4o-mini Stacked Objects Accuracy (%) 87.67 / 97.00 F1-Score 0.9326 / 0.9782 Dynamic Scene Accuracy (%) 94.67 / 98.67 F1-Score 0.9403 / 0.9785
Given the ability of foundational models to interpret the overall context of scenes effectively, we explore whether they can reliably select the appropriate PCB when provided with a question-image pair. We utilize a multi-agent triage model inspired by OpenAI’s Swarm architecture (Bigio et al.,, 2024). As illustrated in Figure 5, our multi-agent framework consists of:
  1. 1.
    A Triage Agent that analyzes the user query and visual input to identify the required type of physical reasoning.
  2. 2.
    Multiple PCBs, each specialized for different physical phenomena (e.g., stability analysis, collision detection, motion tracking).
  3. 3.
    A Foundation Model that receives the PCB-generated context and the original query to produce the response.
In our evaluation, each input consists of a natural language question from a QA dataset paired with a corresponding scene image. This input is initially processed by GPT-4o or GPT-4o-mini that routes the query to one of two specialized PCBs: the PCB for the Falling Tower dataset, or PCB designed for analyzing motion and object interactions as in the CLEVRER dataset. As shown in Table 6, both GPT-4o and GPT-4o-mini achieve excellent F1-Scores, effectively routing queries to the appropriate PCB. These results suggest that PCBs in a multi-agent framework offer a promising approach, with foundation models capable of reliably selecting the correct PCB based on the question-image pair.
5 Discussion
Our results demonstrate that Physics Context Builders (PCBs) offer a promising approach to enhancing physical reasoning in Vision-Language Models. PCBs increase the average performance of GPT-4o (GPT-4o-mini) by 11.1% (11.8%) on CLEVRER and 8.2% (13.8%) on Falling Tower. These improvements are particularly notable given that no modification to the foundation models was required. Despite these gains, large VLMs still exhibit sub-optimal performance in tasks requiring deeper reasoning, such as counterfactual questions (with improvements limited to 1.7-9.5%) and explanatory questions; while enhancing visual perception through PCBs is important, especially for descriptive tasks, the subpar performance on more involved reasoning tasks calls for additional interventions to reach a comprehensive physical understanding.
Besides the performance gain, our findings confirm the value of simulation data in addressing the limitations of VLMs. Unlike simulation-in-the-loop approaches (Wu et al.,, 2015, 2017; Liu et al.,, 2022; Ding et al., 2021b, ) that require computationally expensive simulators during inference, our method uses simulation only to generate synthetic data (context descriptions) for fine-tuning smaller VLMs. This approach is more efficient at inference while still leveraging the rich annotations that simulations can provide. The effectiveness of PCBs in generalizing from simulated to real-world data further supports this approach, as demonstrated by the successful Sim2Real transfer in the Falling Tower benchmark.
PCBs provide a modular, efficient framework for enhancing the physical reasoning capabilities of foundation models without requiring direct fine-tuning. Generating rich physical context from visual inputs uplifts the perceptual ability of foundation models, resulting in more accurate reasoning across diverse physical tasks, from static stability to dynamic collision detection. The multi-agent framework further enhances this approach by enabling adaptive selection of specialized PCBs based on the specific reasoning task.
6 Limitations and Future Work
While our work demonstrates the effectiveness of PCBs across the benchmarks, several limitations remain. First, our current benchmarks focus on a relatively constrained set of physical phenomena—primarily rigid body dynamics and stability. This limits our ability to evaluate how well these approaches might generalize to the full spectrum of physical reasoning that humans perform intuitively, such as reasoning about fluid dynamics or object manipulation.
Second, our framework requires annotated data, which naturally comes from simulation. For unannotated real-world videos, such as those available on YouTube, the lack of structured annotations presents a challenge. An important open problem is how PCBs could be used to extract detailed physical descriptions from real-world videos without explicit annotations, potentially through self-supervised learning or by leveraging other foundation models to generate pseudo-annotations for real data.
Third, the performance improvements, while substantial, still leave room for further enhancement, particularly for complex reasoning tasks like counterfactual/predictive questions. Since PCBs inherently perform direct translation from visual signals to text, they result in strong descriptive capabilities of the foundation models. However, this translation process does not capture all the visual cues needed for more complex reasoning. Hence, we hypothesize that more comprehensive textual descriptions, possibly containing counterfactual/predictive information, could significantly improve the performance of PCBs on these challenging tasks.
A promising future direction is to tackle more complex physical reasoning tasks that require larger-scale simulations. This might involve integrating advanced physics engines that simulate phenomena like fluid dynamics, deformable materials, and articulated mechanisms. Notably, while these simulations could become computationally expensive for simulation-in-the-loop approaches during inference, our PCB framework uses simulation data only for fine-tuning and therefore would not add significant computational burden during inference.
Another avenue for future work is exploring how multiple PCBs could be composed or chained together to handle scenarios requiring reasoning about multiple physical phenomena simultaneously. For instance, reasoning about a scene involving both rigid body dynamics and fluid interactions might benefit from specialized PCBs for each phenomenon, with their outputs combined to provide comprehensive context to the foundation model.
Appendix A Training Details
Here, we provide the training details, including the hyper-parameters, for both QA fine-tuning and PCB training tasks. All of our training was performed on Nvidia’s A100 DGX systems.
A.1 QA fine-tuning
To fine-tune the PaliGemma-3B on the question-answer datasets, we apply the LoRA fine-tuning scheme by targeting the attention weights in both vision and language modules, as well as the fully-connected MLP layers, the multi-modal projector layers, the embedding tokens, the patch embedding, and the positional embedding. Table 7 (left) shows other hyper-parameters used for QA fine-tuning.
Table 7: Hyper-parameters used to fine-tune the PaliGemma-3B models (left) on the QA datasets and (right) as PCB modules.
Hyper-parameter Falling Tower CLEVRER
LoRA rank 16 16
Learning rate 5e-5 5e-5
Batch size 32 32
Epochs 10 3
Trainable parameters 1.24 1.24 %
Number of frames 1 8
[…]
Hyper-parameter Falling Tower CLEVRER
LoRA rank 16 16
Learning rate 5e-5 5e-5
Batch size 32 64
Epochs 50 10
Trainable parameters 1.24 1.24 %
Number of frames 1 8
[…]
A.2 PCB training
Descriptions used for training PCBs. We first discuss the two types of descriptions we considered for training PCBs:
  1. 1.
    Human-Narration (HN), which generates a summary for all the collisions happened in the scene.
    Scene History:
    In this scene, there are 3 collisions occurring in sequence.
    Here are the relevant observations prior to the 1st collision:
    Object 0 (the blue rubber sphere) enters the scene and moves toward the 1st collision site.
    Object 1 (the gray rubber sphere) is moving toward the 1st collision site.
    Object 2 (the cyan metal cube) enters the scene and is moving in the rest of the scene but does not participate in the collision.
    Object 3 (the purple rubber sphere) remains stationary in the scene and does not participate in the collision.
    Object 4 (the blue metal sphere) remains stationary in the scene and does not participate in the collision.
    Finally, Object 0 collides with Object 1.
    Here are the relevant observations prior to the 2nd collision:
  2. 2.
    Structured-Physics (SP), which describes each provided video frame separately as follows, while adding physical properties of the objects, including their discretized and normalized locations and velocities. We also include the location of the collisions happened up to a certain frame.
    [FRAME] [OBJECTS] [OBJ] SHAPE COLOR MATERIAL [LOC] LOC [/LOC] [VEL] VEL [/VEL] [/OBJ] [OBJ] SHAPE COLOR MATERIAL [LOC] LOC [/LOC] [VEL] VEL [/VEL] [/OBJ] ... [/OBJECTS] [COLLISION_PAST] [COLLISION] [OBJ] SHAPE COLOR MATERIAL [/OBJ] [OBJ] SHAPE COLOR MATERIAL [/OBJ] [LOC] LOC [/LOC] [/COLLISION] ... [/COLLISION_PAST] [/FRAME]
Training details. We use the pre-trained PaliGemma-3B model for training the PCB modules and apply the LoRA fine-tuning scheme, similar to the approach used in the QA fine-tuning. Table 7 (right) provides the hyper-parameters used to train PCB modules for both Falling Tower and CLEVRER datasets.
Appendix B Ablations
B.1 The Effect of Framing Multi-Choice Questions as Multiple Binary Questions
As discussed in the main paper, framing the multi-choice questions as multiple binary questions in CLEVRER can yield significant improvement in the accuracy of the models. In Table 8, we provide a comparison between the performance of fine-tuned PaliGemma-3B models with and without this change. As demonstrated, we observe improvement in almost all categories, except for the per question predictive accuracy. We posit that this is because the predictive questions in CLEVRER are always binary questions with exactly one correct choice. Framing the predictive questions as two independent binary questions can result in a model choosing both options as correct or wrong.
Table 8: The performance of the fine-tuned PaliGemma-3B model on question answer pairs for the CLEVRER benchmark based on framing the multi-choice questions as binary questions. Both models are trained for three epochs.
Multi-Choice as Binary? Descriptive Explanatory Predictive Counterfactual
per ques. per opt. per ques. per opt. per ques. per opt.
False 89.3 69.0 86.6 83.6 83.6 41.0 74.0
True 92.9 94.7 98.2 77.9 88.2 68.4 88.7
B.2 The Effect of Training Epochs
We run an ablation study to assess the effect of training for smaller vs. larger number of epochs on the accuracy of CLEVRER in the QA fine-tuning task. Table 9 demonstrates a large improvement in training for more epochs.
Table 9: The performance of the fine-tuned PaliGemma-3B model on question answer pairs for the CLEVRER benchmark based on the number of trained epochs. Here, multi-choise questions are asked as they are (without framing them as multiple binary questions).
Epochs Descriptive Explanatory Predictive Counterfactual
per ques. per opt. per ques. per opt. per ques. per opt.
3 89.3 69.0 86.6 83.6 83.6 41.0 74.0
2 87.2 66.5 85.7 82.3 82.3 38.3 72.9
1 78.1 52.9 77.3 73.5 73.5 15.3 51.0
B.3 Evaluating the Importance of Vision Module
We illustrate the importance of the vision module in a VLM for physical reasoning by conducting the following experiment. Here, we QA-fine-tune only the language model part of PaliGemma-3B while freezing the vision module. The results in Table 10 shows that the performance across all categories drops slightly for the language model-only setting. Therefore, jointly fine-tuning both the vision and language modules is essential for optimal performance, as it enables the model to better align visual features with linguistic representations.
Table 10: Performance drop due to freezing the vision module on the PaliGemma3b-base model for the QA fine-tuning over CLEVRER.
Descriptive Explanatory Predictive Counterfactual
per ques. per opt. per ques. per opt. per ques. per opt.
-1.6 -1.8 -0.6 -8.9 -1.0 -2.8 -1.2
Appendix C Falling Tower Dataset
The Falling Tower dataset is a benchmark for stability detection of stacked objects, inspired by the ShapeStacks benchmark Groth et al., (2018). It includes 4864 unique scenes, 72,775 questions, and detailed simulation-generated annotations to support training Vision-Language Models (VLMs) for spatial and physical reasoning. Each simulation instance is represented as a JSON file containing:
  • Scene Description: A list of objects stacked from bottom to top with their respective offsets, e.g., “Scene description: Here are the parts stacked from bottom to top: purple cube, yellow cylinder. Offsets for each part, from bottom to top, are: (-0.03, -0.05), (0.0, 0.02).”
  • Simulation Metadata: Physical and rendering settings, including stability status (stable: true/false), the number of objects, gravity parameters, and camera settings.
  • Objects: Detailed information about each object, including its type (e.g., cube, cylinder), dimensions, colors (both RGBA and HEX), rigid body properties (e.g., mass, friction), initial and final positions, and positional offsets. Rigid body properties used for simulation were fine-tuned to reflect real-world dynamics, enabling us to achieve 89% accuracy in a human evaluation of 50 examples for stability detection.
  • Questions and Answers: A variety of descriptive and stability QAs aimed at assessing spatial and physical reasoning, e.g:
    • Descriptive Questions: “How many objects are in the scene?” (Answer: 2), “What is the shape/color of the object above the purple cube?” (Answer: yellow cylinder).
    • Stability Questions: “Will this collection of objects stay stationary?” (Answer: False), “Will the yellow cylinder stay stationary?” (Answer: False).
Figure 6 shows the distribution of object stacks in terms of their stability, as well as the distribution of question types.
Figure 6: Falling Tower dataset: (a) distribution of stacked objects in terms of their stability, and (b) distribution of question types, including both descriptive and stability categories.
The Sim2Real dataset consists of 20 images. Seven stable and seven unstable cases were captured against a clean background, while six additional stable cases were captured with a varying background for testing the robustness of a vision model. Additionally, the dataset includes 100 human-generated questions, with five questions per image. The objects are 3D-printed using a J55™ Prime 3D Printer.
Dataset Links: